OpenStudy (anonymous):

I am in need of help again. (t/8)-(8/t)

OpenStudy (lalaly):

\[\frac{t}{8}-\frac{8}{t} = \frac{t}{8} \times \frac{t}{t} -\frac{8}{t} \times \frac{8}{8}\]\[= \frac{t^2-64}{8t}\]

OpenStudy (anonymous):

oh and the equation =0 if that makes a difference. I really hate math lol

OpenStudy (anonymous):



OpenStudy (lalaly):

u mean \[\frac{t^2-64}{8t} =0\]\[=> t^2-64=0 => (t-8)(t+8) =0\]

OpenStudy (anonymous):

(t/8)-(8/t)=0

OpenStudy (lalaly):

yea the equation i wrote is the same as the one u hav,

OpenStudy (lalaly):

but i used a common denominator

OpenStudy (anonymous):

so would the answer be 8 and -8 since it is set to 0? or would it be the simplified fraction?

OpenStudy (lalaly):

yess 8 and-8

OpenStudy (anonymous):

great, thank you for all your help! Glad I found this site today!

OpenStudy (lalaly):

your welcome=)